OpenStudy (anonymous):

Fran Massey earns three times as much as an actuary then she does as a writer. her total income is $40000 more than that of her brother who earns only half as much as Fran does an actuary. What is Frans salary as an actuary?

OpenStudy (anonymous):

She earns x as a writer and 3x as an actuary. When you add that you get 4x = 40000 + A, where A is her brother's salary, which is (3x)/2. When equation 4x=40 000 + (3x)/2 is solved by x, you get x=16 000. So her salary as an actuary is 3 * 16 000 = 48 000.

OpenStudy (anonymous):

Thank you!!!!!!!!!!!